help....use foil
115.) (2-3x6)²


simplify
142.)12) (x³-y³) - (-6x³+x²y-xy²+2y³)



Answer by stanbon(75887)

use foil
115.) (2-3x^6)²
(2-3x^6)(2-3x^6)
F: 2*2 = 4
O: 2*-3x^6 = -6x^6
I: -3x^6*2 = -6x^6
L: -3x^6*-3x^6 = 9x^12
-------------
Combine those products to get:
4 -12x^6 + 9x^12
===============================
simplify
142.)12) (x³-y³) - (-6x³+x²y-xy²+2y³)
= x^2-y^2 + 6x^3 -x^2y+xy^2-2y^3
There are no "like terms to combine.
=====================================
Cheers,
Stan H.